Please point out whether each of the following statements is correct or not.

Carefully explain your reasoning!

a) In Aspartic Acid, the carboxylic group bound to the α carbon has more acidic properties than the carboxylic group that is located in the side chain.

b) The pI of Ala-Gly-Met is 6.0.

c) The peptide Ile-Ala-Pro is more hydrophilic than the peptide Gln-His-Ala.

d) The side chain of Amino Acids that make up a protein function as a buffer.

thank you :)
Answer #1

a) Correct

Explanation: The pKa value of the side-chain COOH = 3.9, that of α-COOH = 3.5

Lesser the pKa, the greater is the acidic nature because of pKa = -Log(Ka)

b) Correct.

Explanation: pI = (3.5 + 9)/2 = 6.25

c) Incorrect.

Explanation: The peptide Gln-His-Ala is more hydrophilic than the peptide Ile-Ala-Pro.